OpenStudy (anonymous):

6.03] Auditions for a play in New York were held with 120 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many men auditioned? 44 40 76 80

OpenStudy (anonymous):

76

OpenStudy (anonymous):

Let x = number of men Therefore (120 - x) = number of women 2x - 3(120 - x) = 20 2x - 360 + 3x = 20 5x = 380 x = 76 (76 men auditioned) 120 - 76 = 44 (44 women auditioned)